OpenStudy (anonymous):

Write the equation for the area (A) in terms of width (w) of a playground if the length (l) is twice as long as the width (w). You do not need to solve the equation.

OpenStudy (anonymous):

do you know the formula for Area (of a rectangle, i presume)?

OpenStudy (anonymous):

No, I don't its been a long time, trying to teach myself all over again to go back to college

OpenStudy (anonymous):

l * w= area of a rectangle

OpenStudy (anonymous):

in this case l=2w so A= 2w*w which equals 2w^2

OpenStudy (anonymous):

can you tell me is this looks right and the answer. A = WL & 2W A = 2W^2 & W = (A/2)^2 positive values only

OpenStudy (anonymous):

are you trying to find A in terms of w? or w in terms of A?

OpenStudy (anonymous):

I don't understand what you are trying to tell me.

OpenStudy (anonymous):

I'm saying that according to the question you asked the answer should be A=2W^2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome